an antique jar is 45 years older than the mug. in 13 years, the jar will be 4 times as old as the mug will be then. find the present age of the jar

Answer by stanbon(75887)

an antique jar is 45 years older than the mug. in 13 years, the jar will be 4 times as old as the mug will be then. find the present age of the jar
--------------
Equations:
j = m + 45
j+13 = 4(m+13)
-----
Substitute for "j" and solve for "m":
m+45 + 13 = 4m+52
3m = 58-52
m = 6/3
m = 2 years (age of the mug)
----
Solve for "j":
j = m + 45
m = 2 + 45
m = 47 years (age of the jar)
==============
Cheers,
Stan H.
===============